OpenStudy (anonymous):

The distance from the sun to the Andromeda galaxy is 1.2×10^19 miles. Light travels at a speed of 5.88×10^12 miles per year (called light year) . How long does it take light to travel from the sun to the Andromeda galaxy ?

OpenStudy (welshfella):

you need to divide 1.2 * 10^19 by 5.88 * 10^12

OpenStudy (anonymous):

@Welshfella isn't the 5.88 supposed to go first ?

OpenStudy (welshfella):

No do this in 2 parts 1/2 / 5.88 = ? 10^19 / 10^12 = ? can you figure these out?

OpenStudy (anonymous):

When I did 1.2 ÷ 5.88 I got 40 but on the calculator it says 0.20408163

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

yes thats correct but i think we can call that 0.204

OpenStudy (welshfella):

now what do you get for 10^19 / 10^12 - recall the law of exponents - when dividing you subtract the exponents

OpenStudy (welshfella):

like x^4 / x^2 = x^(4-2) = x^2

OpenStudy (anonymous):

Yeah so my final answer is 0.204×10^7?

OpenStudy (welshfella):

No i'm afraid not I expect they want the answer in scientific form where the first number must be betwen 1 and 10

OpenStudy (welshfella):

so , to retain the correct value m you multiply the 0.204 by 10 and divide the 10^7 by 10

OpenStudy (anonymous):

Ahh I'm a little confused here

OpenStudy (welshfella):

0.204 * 10 = ? 10^7 / 10^1 = ?

OpenStudy (welshfella):

0.204 * 10^7 is correct but they want it in scientific form

OpenStudy (anonymous):

Kk I got it but since it's a word problem my answer would be "it takes 0.204 miles for light to travel from the sun to Andromeda " m?

OpenStudy (welshfella):

0.204 * 10^7 years

OpenStudy (welshfella):

in scientific notation that is 2.04 * 10^6 years

OpenStudy (anonymous):

It doesn't say to write it in scientific notation

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

yes OK But you should be clear what scientific notation is.

OpenStudy (welshfella):

Notice that the numbers in the question are all in scientific notation - 1.2 and 5.88 are both in the rang 1 - 10.